The demand for tickets to an ethioian camparada film is given by D(p)=200,000-10,000p,where p is price of tickets. If the price of tickets is 12,calculate price elasticity of demand for tickets and draw demand curve

1
Expert's answer
2021-11-14T17:40:53-0500

Given :


D(p)=200,000-10,000 p


At price 12 Birr change demand units will be;

"\frac{Change in demand}{ Change in price }=-10,000"


"D(p)= 200,000 - 10,000\times12"


= 80,000 units


Elasticity in demand= "\frac{Change in demand}{ Change in price unit}"


= "-10,000\times\frac{12}{80,000}"


=-1.5


Demand curve;